# From percolation transition to Anderson localization in one-dimensional speckle potentials

### v1

@@ Abstract
Classical particles in random potentials typically experience a percolation phase transition, being trapped in clusters of mean size $\chi$ that diverges algebraically at a percolation threshold. In contrast, quantum transport in random potentials is controlled by the Anderson localization length, which shows no distinct feature at this classical critical point. Here, we present a comprehensive theoretical analysis of the semi-classical crossover between these two regimes by studying particle propagation in a one-dimensional, red speckle potential, which hosts a percolation transition at its upper bound. As the system deviates from the classical limit, we find that the algebraic divergence of $\chi$ continuously connects to a smooth yet non-analytic increase of the localization length. We characterize this behavior both numerically and theoretically using a semi-classical approach. In this crossover regime, the correlated and non-Gaussian nature of the speckle potential becomes essential, causing the standard DMPK description for uncorrelated disorder to break down. Instead, we predict the emergence of a bimodal transmission distribution, a behavior normally absent in one dimension, which we capture within our semi-classical analysis. Deep in the quantum regime, the DMPK framework is recovered and the universal features of Anderson localization reappear.
@@ 1 Introduction
In quantum disordered systems, subtle interference effects arising from multiple scattering events on random potential fluctuations are responsible for Anderson localization [1, 2]. Over the past decades, this phenomenon has attracted renewed attention thanks to major advances in quantum simulators — most notably ultracold atomic gases, precisely controlled using laser light and other external fields employed to confine the atoms or to tailor various types of disordered potentials [3]. This high level of control has made quantum gases an exceptionally versatile platform for exploring Anderson physics, from weakly interacting systems [4, 5], to setups involving artificial gauge fields [6, 7] or correlated regimes where many-body localization emerges [8]. In dilute gases, these developments have led to seminal observations of Anderson localization in one-dimensional (1D) [9, 10] and three-dimensional [11, 12] configurations, along with detailed investigations of its quantum-chaotic analogue in kicked-rotor models [13, 14].
@@
A distinctive feature of experiments on disordered quantum gases is the use of random potentials exhibiting significant spatial correlations and, in general, non-Gaussian statistics. Among these, optical speckle potentials have been widely employed, and their specific statistical properties have been shown to give rise to peculiar features of Anderson localization, such as effective mobility edges in one dimension [15, 16] or non-monotonic energy dependence of the localization length [17]. Accounting for disorder correlations has also proven crucial for accurately characterizing the localization dynamics of cold atoms in position [18, 9, 19] and momentum [20, 21, 22, 23] space, as well as for proper estimations of the localization length [18, 24], and of the mobility edge in three dimensions [25, 26, 27]. More broadly, correlations may profoundly modify the localization scenario, as exemplified by kicked-rotor systems where they give rise to a diffusive regime in one dimension [28, 13], or induce genuine mobility edges in properly designed quasi-periodic systems (of infinite-range correlations) [29, 30, 31]. In the context of optics, light in correlated media has also been shown to exhibit a variety of transport regimes [32].
@@
Another remarkable feature of speckle potentials is their ability to support a percolation transition. In classical physics, percolation theory investigates the conditions under which randomly distributed objects form connected structures (“clusters”) that enable transport over arbitrarily long distances [33, 34, 35]. The critical point at which such large-scale connectivity emerges is known as the percolation threshold; it marks a geometric phase transition characterized by nontrivial critical behavior. In continuous speckle potentials, percolation thresholds have been identified in one, two, and three dimensions, corresponding to the energy level above which a classical particle can travel infinitely far along potential valleys [36, 11, 37]. At the quantum level, however, the transport scenario changes drastically. In one and two dimensions, quantum particles are always localized, whereas in three dimensions an Anderson metal–insulator transition occurs [38]—but at a point that differs markedly from the classical percolation threshold [11, 39]. Quantum mechanically, no direct trace of a critical percolation transition therefore appears to survive. Still, as one approaches the classical limit the Anderson and percolation problems should coincide, and the percolation transition re-emerge. How this connection establishes is the question we address in the present work, focusing on particle transport in a 1D speckle disorder—one of the simplest yet nontrivial systems exhibiting a percolation transition.
@@
In detail, in this work we theoretically characterize the localization length near the percolation threshold as the system approaches its classical limit. This limit is governed by an effective Planck constant, $\hbar_{\mathrm{eff}}$, measuring the ratio between the de Broglie wavelength and the disorder correlation length. As this ratio increases from zero, we find that the percolation transition smoothens, with the localization length crossing over from a strict divergence at the percolation threshold when $\hbar_{\mathrm{eff}}=0$, to a continuous—though extremely steep— growth as $\hbar_{\mathrm{eff}}\ll 1$ becomes nonzero. We describe this crossover both numerically and theoretically, by combining advanced properties of speckle statistics with semi-classical tools [40, 41]. By analyzing the localization length and the statistical distribution of the transmission, we further show that near the percolation threshold the standard 1D Anderson localization theory for uncorrelated Gaussian disorder (Fokker-Planck/DMPK equation) [42, 43] fails completely. In particular, we find that the transmission distribution exhibits a bimodal structure typical of diffusive transport, usually absent in 1D models of uncorrelated disorder. The characteristic universal features of Anderson localization re-emerge only for larger $\hbar_{\text{eff}}$, in particular in the deep quantum limit $1/\hbar_{\text{eff}}\ll 1$ where usual, quantum-mechanical perturbative expansions at weak disorder can be applied [15].
@@
The article is organized as follows. Section 2 recalls the main properties of the red-detuned random potential studied throughout this work. We define the semi-classical transport regime and derive the critical, classical scaling of the localization length at the percolation threshold. In Sec. 3, we first attempt to connect Anderson localization theory to this classical limit by defining the localization length from the logarithm of individual speckle potential barriers. We show that this approach is insufficient and that accessing the classical regime requires a more appropriate definition. A semi-classical treatment valid for $\hbar_{\text{eff}}\ll 1$ and resolving these limitations is then presented in Sec. 4, and benchmarked against exact transfer-matrix simulations. In Sec. 5, we extend the analysis to the full statistical distribution of the particle transmission across the percolation threshold. In the subcritical region, this distribution becomes bimodal — a feature not captured by the standard 1D Fokker-Planck equation but well reproduced by our semi-classical approach. Finally, Sec. 6 discusses the crossover to the deep quantum regime ($\hbar_{\text{eff}}\gg 1$), where perturbation theory in $1/\hbar_{\text{eff}}$ becomes applicable.
@@ 2.1 Speckle disorder and semi-classical regime
where $\sigma$ denotes the correlation length. A typical realization of such a random potential is shown in Fig. 1(a). In practice, red- or blue-detuned correlated speckle potentials are commonly produced in cold-atom experiments by shining the atoms with a laser reflected or transmitted through a rough surface. Attractive (resp. repulsive) potentials experienced by the atoms are obtained by tuning the laser far to the red (resp. blue) of a relevant atomic transition [45, 5].
Figure 1: (a) The model: we consider the transmission of a quantum particle of energy $E\sim V_{0}\gg E_{\sigma}$ through a 1D red-detuned speckle potential lying between $x=0$ and $x=L$. The potential $V(x)<V_{0}$ is bounded from above and has a correlation length $\sigma$. (b) For a large enough system size, a classical particle with energy $E<V_{0}$ always ends trapped within a potential valley (cluster). These clusters correspond to regions of space where EV(x)>0E-V(x)>0, and have a characteristic size $\chi$. Near $E=V_{0}$, the speckle potential looks like a succession of inverted harmonic bumps of mean size $\sigma$.
@@
This regime corrresponds to a physical situation where the particle flies near the classical percolation threshold, and has a de Broglie wavelength $h/\sqrt{2mE}$ much smaller than the disorder correlation length $\sigma$. Under these conditions, the particle effectively resolves all potential fluctuations and experiences negligible tunneling, thereby behaving almost classically. The condition (3) introduces the relevant dimensionless parameter
@@ 2.2 Percolation transition and critical scaling of transmission
In the strict classical limit $\hbar_{\text{eff}}=0$, the percolation threshold at $E=V_{0}$ corresponds to a phase transition that manifests itself through the critical behavior of a characteristic cluster size. In this work, we are primarily interested in how this critical phenomenon appears in particle transport, which can be naturally characterized by the average transmission coefficient $\langle T\rangle$ through the disordered region, where the brackets denote averaging over the speckle statistics. Near the percolation threshold, its classical value $\langle T_{\text{class}}\rangle$ is expected to take the form $\langle T_{\text{class}}\rangle\sim\exp\left[-{L}/{\chi(E)}\right]$, where $\chi(E)$ is a characteristic cluster size [see Fig. 1(b)] behaving as
@@
with $\gamma$ a critical exponent. In terms of the topographic properties of the speckle potential, the factor $\exp\left(-{L}/{\chi}\right)$ can alternatively be interpreted as the probability distribution of clusters of size $L$: this distribution is essentially nonzero for $L<\chi(E)$, indicating that at a given energy $E$, most clusters have sizes smaller than $\chi(E)$.
@@
$P(V,\omega)\simeq\frac{2\sqrt{2}}{\omega_{0}}\frac{1}{\sqrt{\pi(V_{0}-V)V_{0}}}\left(\frac{\omega}{\omega_{0}}\right)^{2}\exp\left(-\frac{\omega^{2}}{\omega_{0}^{2}}\right),$ (6)
@@
The distribution $P(V,\omega)$ can be used to compute $\langle T_{\text{class}}\rangle$, by noticing that near $E=V_{0}$ the classical transmission is the product of the average transmissions through uncorrelated inverted harmonic barriers, each of typical size $\sigma$. Denoting by $T_{b}$ the transmission coefficient through a single barrier of maximum $V$, we have $T_{b}=\Theta(E-V)$ since a classical particle is either transmitted or reflected depending on whether its energy $E$ is larger of smaller than $V$. The total transmission is therefore
@@
$\langle T_{\text{class}}\rangle=\exp\left[-\frac{L}{\chi(E)}\right],\ \ \ \chi(E)\underset{E\to V_{0}^{-}}{\simeq}\frac{1}{\rho\sqrt{2(V_{0}-E)/V_{0}}}.$ (10)
@@ 3 Localization length from logarithmic transmission
$\xi^{-1}(E)=-\rho\,\langle\text{ln}\,T_{j}\rangle$ (11)
defined from the logarithm of the transmission $T_{j}$ through a single potential barrier. A quick look at this formula highlights an obvious difficulty arising for energies $E<V_{0}$, if one tries to evaluate it in the classical limit $\hbar_{\text{eff}}=0$:
@@
$T_{b}=\frac{1}{1+\exp\left[-\frac{2\pi}{\hbar\omega}(E-V)\right]}.$ (13)
@@
$\xi^{-1}(E)=\rho\int_{0}^{\infty}d\omega\int_{-\infty}^{V_{0}}dV\,P(V,\omega)\,\text{ln}\left(1+\exp\left[-\frac{2\pi}{\hbar\omega}(E-V)\right]\right).$ (14)
@@
$\xi^{-1}(E)\simeq\rho\int_{0}^{\infty}d\omega\left\{\int_{-\infty}^{E}dV\,\exp\left[-\frac{2\pi}{\hbar\omega}(E-V)\right]+\int_{E}^{V_{0}}dV\frac{2\pi}{\hbar\omega}(V-E)\right\}P(V,\omega).$ (15)
@@
Both the classical result (3) or its semi-classical version (16) thus confirm that the localization length defined according to (11) does not converge to $\chi$ [Eq. (10)], which is the proper critical behavior expected at the percolation threshold. The reason for this mismatch is evident from the above calculation: the localization length defined via (11) describes a propagation dominated by tunneling through barriers, while completely neglecting the cluster contributions $E>V$. This is legitimate in the deep quantum regime $\hbar_{\text{eff}}\gg 1$ (see Sec. 6.2), but is clearly inadequate in the semi-classical regime (3) where the particle energy exceeds the height of most encountered barriers. From a mathematical standpoint, the reason why Eq. (11) incorrectly captures the decay of $\langle T\rangle$ when $\hbar_{\text{eff}}\ll 1$ lies in the fact that it originates from a model in which disorder is assumed to be both uncorrelated and Gaussian distributed [43, 42]. These assumptions are fundamentally incompatible with the semi-classical condition (3), for two reasons. First, because $\hbar_{\text{eff}}$ diverges as the disorder correlation length $\sigma\to 0$, making the classical condition $E\gg E_{\sigma}$ never fulfilled. Second, because near the upper bound $E\sim V_{0}$, the non-Gaussian character of the red-speckle potential becomes most pronounced, while a Gaussian model would instead describe a potential with arbitrarily high peaks, obviously not bounded. In the next section, we present a correct determination of $\langle T\rangle$ and of the localization length in the semi-classical limit.

### v2

@@ Abstract
Classical particles in random potentials typically experience a percolation phase transition, being trapped in clusters of mean size $\chi$ that diverges algebraically at a percolation threshold. In contrast, quantum transport in random potentials is controlled by the Anderson localization length, which shows no distinct feature at this classical critical point. Here, we present a comprehensive theoretical analysis of the semi-classical crossover between these two regimes by studying particle propagation in a one-dimensional, red speckle potential, which hosts a percolation transition at its upper bound. As the system deviates from the classical limit, we find that the algebraic divergence of $\chi$ continuously connects to a smooth yet non-analytic increase of the localization length. We characterize this behavior both numerically and theoretically using a semi-classical approach. In this crossover regime, the correlated and non-Gaussian nature of the speckle potential becomes essential, causing the standard Dorokhov-Mello-Pereyra-Kumar (DPMK) description for uncorrelated disorder to break down. Instead, we predict the emergence of a bimodal transmission distribution, a behavior normally absent in one dimension, which we capture within our semi-classical analysis. Deep in the quantum regime, the DMPK framework is recovered and the universal features of Anderson localization reappear.
@@ 1 Introduction
In quantum disordered systems, subtle interference effects arising from multiple scattering events on random potential fluctuations are responsible for Anderson localization [1, 2]. Over the past decades, this phenomenon has attracted renewed attention thanks to major advances in quantum simulators—most notably ultracold atomic gases, precisely controlled using laser light and other external fields employed to confine the atoms or to tailor various types of disordered potentials [3]. This high level of control has made quantum gases an exceptionally versatile platform for exploring Anderson physics, from weakly interacting systems [4, 5], to setups involving artificial gauge fields [6, 7] or correlated regimes where many-body localization emerges [8]. In dilute gases, these developments have led to seminal experimental observations of Anderson localization in one-dimensional (1D) [9, 10] and three-dimensional [11, 12] configurations, along with detailed investigations of its quantum-chaotic analogue in kicked-rotor models [13, 14].
@@
A distinctive feature of experiments on disordered quantum gases is the use of random potentials exhibiting significant spatial correlations and, in general, non-Gaussian statistics. Among these, optical speckle potentials have been widely employed, and their specific statistical properties have been shown to give rise to peculiar features of Anderson localization, such as effective mobility edges in one dimension [15, 16] or non-monotonic energy dependence of the localization length [17]. Accounting for disorder correlations has also proven crucial for accurately characterizing the localization dynamics of cold atoms in position [18, 9, 19] and momentum [20, 21, 22, 23] space, as well as for proper estimations of the localization length [18, 24], and of the mobility edge in three dimensions [25, 26, 27]. More broadly, correlations may profoundly modify the localization scenario, as exemplified by kicked-rotor systems where they give rise to a diffusive regime in one dimension [28, 13], or induce genuine mobility edges in appropriately designed quasi-periodic systems (of infinite-range correlations) [29, 30, 31]. In the context of optics, light in correlated media has also been shown to exhibit a variety of transport regimes [32].
@@
Another interesting feature of disordered speckle potentials is their percolation transition. In classical physics, percolation theory investigates the conditions under which randomly distributed objects form connected structures (“clusters”) that enable transport over arbitrarily long distances [33, 34, 35]. The critical point at which such large-scale connectivity emerges is known as the percolation threshold; it marks a geometric phase transition characterized by nontrivial critical behavior. In continuous speckle potentials, percolation thresholds have been identified in one, two, and three dimensions, corresponding to the energy level above which a classical particle can travel infinitely far along potential valleys [36, 11, 37]. At the quantum level, however, the transport scenario changes drastically. In one and two dimensions, quantum particles are always localized, whereas in three dimensions an Anderson metal-insulator transition occurs [38]—but at a point that differs markedly from the classical percolation threshold [11, 39]. Quantum mechanically, no direct trace of a critical percolation transition therefore appears to survive. Still, as one approaches the classical limit the Anderson and percolation problems should coincide, and the percolation transition re-emerge. How this connection establishes is the question we address in the present work, focusing on particle transport in a 1D speckle disorder—a simple and experimentally relevant continuous model which exhibits a percolation transition.
@@
In detail, in this work we theoretically characterize the localization length near the percolation threshold as the system approaches its classical limit. This limit is governed by an effective Planck constant, $\hbar_{\mathrm{eff}}$, measuring the ratio between the de Broglie wavelength and the disorder correlation length. As this ratio increases from zero, we find that the percolation transition smoothens, with the localization length crossing over from a strict divergence at the percolation threshold when $\hbar_{\mathrm{eff}}=0$, to a continuous—though extremely steep— growth as $\hbar_{\mathrm{eff}}\ll 1$ becomes nonzero. We describe this crossover both numerically and theoretically, by combining advanced properties of speckle statistics with semi-classical tools [40, 41]. By analyzing the localization length and the statistical distribution of the transmission, we further show that near the percolation threshold the standard 1D Anderson localization theory for uncorrelated Gaussian disorder (Fokker-Planck/Dorokhov-Mello-Pereyra-Kumar or DMPK equation) [42, 43] fails completely. In particular, we find that the transmission distribution exhibits a bimodal structure typical of diffusive transport, usually absent in one dimension. The characteristic universal features of Anderson localization re-emerge only for larger $\hbar_{\text{eff}}$, in particular in the deep quantum limit $1/\hbar_{\text{eff}}\ll 1$ where usual, quantum-mechanical perturbative expansions at weak disorder can be applied [15].
@@
The article is organized as follows. Section 2 recalls the main properties of the red-detuned random potential studied throughout this work. We define the semi-classical transport regime and derive the critical, classical scaling of the localization length at the percolation threshold. In Sec. 3, we first attempt to connect Anderson localization theory to this classical limit by defining the localization length from the logarithm of individual speckle potential barriers. We show that this approach is inadequate and that accessing the classical regime requires a more appropriate definition. A semi-classical treatment valid for $\hbar_{\text{eff}}\ll 1$ and resolving these limitations is then presented in Sec. 4, and benchmarked against exact transfer-matrix simulations. In Sec. 5, we extend the analysis to the full statistical distribution of the particle transmission across the percolation threshold. In the subcritical region, this distribution becomes bimodal—a feature not captured by the standard 1D Fokker-Planck equation but well reproduced by our semi-classical approach. Finally, Sec. 6 discusses the crossover to the deep quantum regime ($\hbar_{\text{eff}}\gg 1$), where perturbation theory in $1/\hbar_{\text{eff}}$ becomes applicable.
@@ 2.1 Speckle disorder and semi-classical regime
where $\sigma$ denotes the correlation length. A typical realization of such a random potential is shown in Fig. 1(a). In practice, correlated speckle potentials are commonly produced in cold-atom experiments by shining the atoms with a laser reflected or transmitted through a rough surface. A red-detuned (resp. blue-detuned) speckle then corresponds to an attractive (resp. repulsive) potential experienced by the atoms, obtained by tuning the laser far to the red (resp. blue) of a relevant atomic transition [45, 5].
Figure 1: (a) The model: we consider the transmission of a quantum particle of energy $E\sim V_{0}\gg E_{\sigma}$ through a 1D red-detuned speckle potential lying between $x=0$ and $x=L$. The potential $V(x)<V_{0}$ is bounded from above and has a correlation length $\sigma$. (b) Near $E=V_{0}$, the speckle potential looks like a succession of inverted harmonic bumps of mean size $\sigma$. For a large enough system size, a classical particle with energy $E<V_{0}$ coming from the left always ends trapped within a potential valley (cluster). These clusters correspond to regions of space where EV(x)>0E-V(x)>0, and have a characteristic size $\chi$.
@@
This regime corresponds to a physical situation where the particle flies near the classical percolation threshold, and has a de Broglie wavelength $h/\sqrt{2mE}$ much smaller than the disorder correlation length $\sigma$. Under these conditions, the particle effectively resolves all potential fluctuations and experiences negligible tunneling, thereby behaving almost classically. The condition (3) introduces the relevant dimensionless parameter
@@ 2.2 Percolation transition and critical scaling of transmission
In the strict classical limit $\hbar_{\text{eff}}=0$, the percolation threshold at $E=V_{0}$ corresponds to a phase transition that manifests itself through the critical behavior of a characteristic cluster size. In this work, we are primarily interested in how this critical phenomenon appears in particle transport, which can be naturally characterized by the average transmission coefficient $\langle{T}\rangle$} through the disordered region, where the brackets denote averaging over the speckle statistics. Near the percolation threshold, its classical value $\langle{T_{\text{class}}}\rangle$} is expected to take the form ${\langle{T_{\text{class}}}\rangle}\sim\exp[-{L}/{\chi(E)}]$, where $\chi(E)$ is a characteristic cluster size [see Fig. 1(b)] behaving as
@@
with $\gamma$ a critical exponent. In terms of the topographic properties of the speckle potential, the factor $\exp(-{L}/{\chi})$ can alternatively be interpreted as the probability distribution of clusters of size $L$: this distribution is essentially nonzero for $L<\chi(E)$, indicating that at a given energy $E$, most clusters have sizes smaller than $\chi(E)$.
@@
$P(V,\omega)\simeq\frac{2\sqrt{2}}{\omega_{0}}\frac{1}{\sqrt{\pi(V_{0}-V)V_{0}}}\left(\frac{\omega}{\omega_{0}}\right)^{2}\exp(-\frac{\omega^{2}}{\omega_{0}^{2}})\,$, (6)
@@
The distribution $P(V,\omega)$ can be used to compute $\langle{T_{\text{class}}}\rangle$}, by noticing that near $E=V_{0}$ the classical transmission is the product of the average transmissions through uncorrelated inverted harmonic barriers, each of typical size $\sigma$. Denoting by $T_{b}$ the transmission coefficient through a single barrier of maximum $V$, we have $T_{b}=\Theta(E-V)$ since a classical particle is either transmitted or reflected depending on whether its energy $E$ is larger or smaller than $V$. The total transmission is therefore
@@
${\langle{T_{\text{class}}}\rangle}=\exp[-\frac{L}{\chi(E)}],\ \ \ \chi(E)\underset{E\to V_{0}^{-}}{\simeq}\frac{1}{\rho\sqrt{2(V_{0}-E)/V_{0}}}.$ (10)
@@ 3 Localization length from logarithmic transmission
$\xi^{-1}(E)=-\rho\,{\langle{\text{ln}\,T_{b}}\rangle}$ (11)
defined from the logarithm of the transmission $T_{b}$ through a single potential barrier. A quick look at this formula highlights an obvious difficulty arising for energies $E<V_{0}$, if one tries to evaluate it in the classical limit $\hbar_{\text{eff}}=0$:
@@
$T_{b}=\frac{1}{1+\exp[-\frac{2\pi}{\hbar\omega}(E-V)]}\,.$ (13)
@@
$\xi^{-1}(E)=\rho\int_{0}^{\infty}d\omega\int_{-\infty}^{V_{0}}dV\,P(V,\omega)\,\ln(1+\exp[-\frac{2\pi}{\hbar\omega}(E-V)]).$ (14)
@@
$\xi^{-1}(E)\simeq\rho\int_{0}^{\infty}d\omega\left\{\int_{-\infty}^{E}dV\,\exp[-\frac{2\pi}{\hbar\omega}(E-V)]+\int_{E}^{V_{0}}dV\frac{2\pi}{\hbar\omega}(V-E)\right\}P(V,\omega)\,.$ (15)
@@
Both the classical result (3) or its semi-classical version (16) thus confirm that the localization length defined according to (11) does not converge to $\chi$ [Eq. (10)], which is the proper critical behavior expected at the percolation threshold. The reason for this mismatch is evident from the above calculation: the localization length defined via (11) describes a propagation dominated by tunneling through barriers, while completely neglecting the cluster contributions $E>V$. This is legitimate in the deep quantum regime $\hbar_{\text{eff}}\gg 1$ (see Sec. 6.2), but is clearly inadequate in the semi-classical regime (3) where the particle energy, on the contrary, exceeds the height of most encountered barriers. From a mathematical standpoint, the reason why Eq. (11) incorrectly captures the decay of $\langle T\rangle$ when $\hbar_{\text{eff}}\ll 1$ lies in the fact that it originates from a model in which disorder is assumed to be both uncorrelated and Gaussian distributed [43, 42]. These assumptions are fundamentally incompatible with the semi-classical condition (3) for two reasons. First, because $\hbar_{\text{eff}}$ diverges as the disorder correlation length $\sigma\to 0$, making the classical condition $E\gg E_{\sigma}$ never fulfilled. Second, because near the upper bound $E\sim V_{0}$, the non-Gaussian character of the red-detuned speckle potential becomes most pronounced, while a Gaussian model would instead describe a potential with arbitrarily high peaks, obviously not bounded. In the next section, we present a correct determination of $\langle{T}\rangle$} and of the localization length in the semi-classical limit.